OpenStudy (anonymous):

Identify the correct slope and y intercept of the equation 2x + 4y = 12. A. slope = 2; y intercept at (0, 12) B.slope = one half; y intercept at (0, 3) C. slope = one half; y intercept at (0, 4) D. slope = -one half; y intercept at (0, 3) i think it's D but im not sure...

OpenStudy (lgbasallote):

\[\implies 4y = -2x + 12\] \[\implies y = \frac{-2x + 12}{4}\] \[\implies y = -\frac 24x + \frac{12}{4}\] \[\implies y = -\frac 12 x + 3\] in short you're right

OpenStudy (anonymous):

ok thank